OpenStudy (pratyush5):

A body looses half of its velocity while penetrating 3 cm into a wooden block, then how much will it penetrate more before finally coming to rest. Please explain

OpenStudy (pratyush5):

@ashwinjohn3 @yrelhan4 @uri

OpenStudy (pratyush5):

i have got the answer but i want to know a proper method

OpenStudy (yrelhan4):

newtons law of motion will do it for you!

OpenStudy (pratyush5):

i used v^2 - u^2 = 2as

OpenStudy (pratyush5):

i just need to know what should i take for acceleration or should i ignore it

OpenStudy (anonymous):

u must find the acceleration dude....

OpenStudy (anonymous):

Ok....Acceleration is negative since it is slowing down.....

OpenStudy (pratyush5):

ya sure

OpenStudy (anonymous):

here v=u/2,and use the equation v^2=u^2-2as

OpenStudy (anonymous):

then u will get accelration

OpenStudy (anonymous):

then finally equate v=0,a=obtained from above and find s!!!

OpenStudy (anonymous):

since here the question asked is "how much more will it penetrate" initial velocity=u/2...

OpenStudy (pratyush5):

i am getting a= -u^2/8

OpenStudy (pratyush5):

is dat right ?

OpenStudy (yrelhan4):

yup..

OpenStudy (anonymous):

yup....

OpenStudy (pratyush5):

there fore the equation for finding "s" should be right